Given a normal distribution, X, with mean, 110, and standard deviation, sigma = 46.

A. What is the X value with Z-score equal to z = 0.65?

B. What is the probability of X is less than or equal to 146.9? %
Answer #1

solution:

Given that,

mean = = 110

standard deviation = = 46

Using standard normal table,

= P(Z < z) = 0.65

= P(Z <0.39 ) = 0.65

z = 0.39 Using standard normal table,

Using z-score formula  

x= z * +

x= 0.39*46+110

x= 127.94

b.

P(X 146.9) = P[(X- ) / (146.9-110) /46 ]

= P(z 0.80)

Using z table

= 0.7881